# A Comprehensive Survey on Machine Learning Driven Material Defect Detection: Challenges, Solutions, and Future Prospects

### v1

@@ Abstract.
Material defects represent a primary challenge affecting product performance and giving rise to safety issues in related products. The rapid and accurate identification and localization of material defects constitute crucial research endeavours in addressing contemporary challenges associated with material defects. Although conventional non-destructive testing methods such as ultrasonic and X-ray approaches have mitigated issues related to low efficiency in manual inspections, they struggle to meet the diverse requirements of high precision, real-time speed, automation, and intelligence. In recent years, propelled by the swift advancement of machine learning technologies, particularly exemplified by deep learning, machine learning has swiftly emerged as the core technology and a prominent research direction for material defect detection. Through a comprehensive review of the latest literature, we systematically survey the machine learning techniques applied in material defect detection into five categories: unsupervised learning, supervised learning, semi-supervised learning, reinforcement learning, and generative learning. We provide a detailed analysis of the main principles and techniques used, together with the advantages and potential challenges associated with these techniques. Furthermore, the survey focuses on the techniques for defect detection in composite materials, which are important types of materials enjoying increasingly wide application in various industries such as aerospace, automotive, construction, and renewable energy. Finally, the survey explores potential future directions in material defect detection utilizing machine learning technologies. This comprehensive survey not only consolidates existing literature on machine learning-based material defect detection technologies but also serves as a foundational reference for future researchers and industrial practitioners, providing valuable insights and guidance in developing advanced and efficient material defect detection systems.
@@
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456conference: ACM Computing Surveys; ACM Computing Surveysbooktitle: ;price: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Material defect detection (MDD) stands as a critical subfield within the broader domain of defect detection, with a primary focus on identifying and assessing defects in various materials. MDD is a pivotal element in ensuring product quality and production efficiency, and it holds particular significance in the manufacturing industry. The defects can occur on the surface or internally, potentially compromising the product’s performance and safety (usamentiaga2022automated, ). As such, MDD is an essential component in numerous industries, playing a vital role in ensuring the safety, reliability, and quality assurance of products. Adherence to strict regulatory standards is crucial, not only to prevent catastrophic failures and reduce waste but also to maintain cost efficiency. Effective defect detection methodologies directly contribute to the longevity and performance of products, thereby driving innovation in manufacturing processes. Additionally, they play a key role in promoting environmental sustainability by minimizing material waste. In this regard, MDD is fundamental to the responsible and successful production and distribution of goods across various sectors.
@@
Current MDD methods must adapt to the increasing complexity of manufacturing processes and product diversity to meet the demands of modern manufacturing. As a result, manufacturers face a series of challenges and higher detection requirements. One major challenge is the identification of complex defects. The growing variety of materials and the complexity of manufacturing processes lead to increasingly intricate defect types within materials, such as cracks, pores, and inclusions. These defects exhibit diverse shapes, sizes, and positions, making their identification difficult. Another challenge is the detection of minute defects. Some material defects are exceedingly small, even imperceptible to the naked eye, and require high-precision detection methods to identify their potentially significant impacts on product performance and safety. Furthermore, MDD often involves handling multimodal data, including images, ultrasonic waves, and X-rays. Each data type has distinct features and noise levels, posing a challenge in effectively processing and integrating these diverse sources of information. Real-time online detection of defects in materials is also crucial for timely issue identification and resolution on production lines (shi2020overview, ; ou2021recent, ). This requirement necessitates the development of methods with fast, stable, and reliable performance. The ability to detect defects in real-time enables manufacturers to promptly address issues, minimize production interruptions, and maintain high product quality standards (yang2020using, ; zeng2022small, ; zhao2023multi, ). Moreover, the rarity of defective samples also poses difficulties in collecting sufficient data for analysis. Therefore, developing effective and efficient MDD techniques that can operate optimally with these practical challenges is becoming increasingly crucial (jin2022survey, ; shahrabadi2022defect, ).
Traditional approaches to MDD often depend on manual inspection, which comes with significant limitations. As highlighted by Bhatt et al. (bhatt2021image, ), manual inspection is subject to the subjective judgment of inspectors and generally lacks precision, particularly when dealing with minuscule defects or in environments that are potentially harmful to humans. Moreover, manual inspection struggles to keep pace with the increasing complexity of products and the accelerating speeds of modern manufacturing processes. Consequently, manual inspection is often insufficient to meet the rigorous demands of contemporary production standards, necessitating the adoption of more advanced and reliable methods for surface defect detection in industrial products.
@@
As a response to these challenges, the implementation of visual perception-based surface defect detection has gained prominence as a more effective alternative, progressively replacing manual inspections. Machine vision and computer vision, as rapidly evolving branches in the field of artificial intelligence, have found extensive applications across various industrial domains, including the critical area of MDD (hafiz2021attention, )(ren2022state, ). This advanced technology, which employs non-contact automatic detection, offers higher accuracy and demonstrates reliable performance in complex production environments. The integration of these technologies into MDD involves the automated inspection and analysis of product surfaces or interiors, utilizing advanced image processing and computer vision algorithms. These cutting-edge technologies enable the automatic identification and classification of defects, proving capable of handling vast volumes of data and conducting high-precision inspections rapidly. Moreover, the application of machine vision and computer vision significantly enhances the accuracy and consistency of defect detection. By reducing the reliance on human inspection, these technologies diminish the impact of human error, contributing to more reliable and uniform inspection results. This improvement in detection quality, in turn, helps to lower production costs and boost overall production efficiency (ren2022state, ; luo2020automated, ; fan2020line, ). The effectiveness of visual perception-based surface defect detection has been proven across various sectors. For instance, Stojanovic et al. (stojanovic2001real, ) showcased its application in the tile industry, while Wang et al. (wang2021defect, ) and Dong et al. (dong2021automatic, ) have made notable innovations in the fabric sector. Similarly, significant contributions to steel plate defect detection have been made by Hu et al. (hu2020unsupervised, ) and Wang et al. (wang2018distributed, ). Furthermore, the technology has been successfully applied in PCB defect detection, as evidenced by the works of Zhang et al. (zhang2021semi, ) and Mei et al. (mei2018unsupervised, ).
Machine learning (ML), a subset of artificial intelligence, has been the driving force behind automated defect detection based on machine vision and computer vision. Since its inception in the mid-20th century, ML has evolved significantly, from early concepts like the Turing Test and the development of learning algorithms to the adoption of statistical methods and the breakthroughs in deep learning (DL) and neural networks (janiesch2021machine, ). The advent of big data and increased computational power has led to a resurgence in the field, with applications spanning various sectors, including healthcare, finance, and autonomous vehicles (cioffi2020artificial, ). ML plays a pivotal role in MDD by automating the entire detection process, from data preprocessing and feature extraction to defect identification and classification. This automation reduces manual intervention, increases efficiency, and lowers detection costs (nturambirwe2020machine, ). Trained ML algorithms exhibit heightened sensitivity, enabling the precise identification of extremely small defects, even those imperceptible to the human eye. This significantly improves the accuracy of MDD and facilitates early issue detection, preventing declines in product quality. Moreover, ML models ensure sustained and stable detection capabilities over extended periods, unaffected by factors like fatigue or emotions, thereby enhancing reliability and consistency. These models excel at processing large volumes of data in short timeframes, efficiently handling defect detection tasks for numerous products or images and meeting the demands of large-scale production. In addition to real-time defect detection, ML technologies provide valuable data support for quality control through offline data analysis. This analysis contributes to process improvements, enhanced product quality, and the discovery of new defect patterns and trends in historical data. By uncovering these insights, ML guides scientific refinements in manufacturing processes, enabling continuous optimization and reducing the occurrence of defects (sresakoolchai2022railway, ; cao2020review, ). The integration of ML in MDD not only streamlines the detection process but also fosters a data-driven approach to quality control, ultimately leading to higher-quality products and more efficient manufacturing processes.
Despite the high detection efficiency and accuracy of ML in MDD, there are several drawbacks to consider. One significant limitation is the poor generality of detection algorithms, as they may not be readily applicable to other materials or contexts, hindering widespread adoption across diverse manufacturing settings. Another drawback is the high requirements for the quality of defect visual images, as complex backgrounds and excessive noise can increase the complexity of detection algorithms and reduce their accuracy. Obtaining high-quality, clean, and well-prepared visual data can be time-consuming and resource-intensive, especially in industrial environments with varying conditions. Furthermore, the maintenance and production costs for machine vision detection systems can be substantial due to the need for specialized hardware, software, and expertise. The initial investment and ongoing costs for upgrades, repairs, and personnel training may be a barrier for some manufacturers, particularly small and medium-sized enterprises with limited resources (cui2021sddnet, ; caiazzo2022towards, ).
With the development of new ML technologies, particularly those centred around DL, there has been an increasing application in MDD. DL, a branch of ML that utilizes artificial neural networks to simulate the operation of the human brain’s neural network, aims to process and analyze complex data (lecun2015deep, ). Common types of deep networks include: (1) Convolutional Neural Networks (CNN), which are used for processing image data, extracting features layer by layer, and demonstrating robust feature learning and classification capabilities, with common models including LeNet, AlexNet, VGGNet, GoogLeNet, and ResNet; (2) Recurrent Neural Networks (RNN), which are applied for handling sequential data such as text, speech, and time series, and can be used for real-time monitoring and prediction of continuous data on production lines in MDD; (3) Long Short-Term Memory Networks (LSTM), a special type of RNN effective in handling long sequential data, addressing the gradient vanishing problem seen in traditional RNNs, and applicable for real-time monitoring and prediction of continuous image or production line data in MDD; and (4) Generative Adversarial Networks (GAN), which comprise a generator and a discriminator that competitively achieve unsupervised learning of data, and can be employed to generate realistic defect samples, enhancing the model’s generalization capability in MDD (shrestha2019review, ).
By constructing suitable network structures, DL can effectively address the issue in traditional machine vision detection algorithms, where different algorithms are required for different MDD tasks. Various DL technologies can extract high-level features from images without the need for manual feature design, facilitating defect detection and recognition based on these features and significantly advancing the field of MDD (hafiz2021attention, ). These technologies can meet the multiple requirements of contemporary MDD, including automation, intelligence, high detection efficiency, high detection accuracy, and real-time detection (bhatt2021image, ). The proliferation of high-performance computers and GPUs has also made the application of DL in MDD more convenient and cost-effective, as DL requires substantial computational resources.
@@
The existing literature on defect detection predominantly focuses on general methodologies or specific application domains (P2, ; P6, ; P7, ; P8, ; P10, ; P11, ), with some studies delving into defect detection in particular types of materials (P3, ; P5, ; P9, ; P17, ). However, a systematic review of ML-driven MDD remains largely unexplored. Our survey presents several distinct advantages that highlight its contributions to the field:
@@ 2. Preliminaries and Problem Identification
In this section, we first introduce the fundamental concepts and procedural aspects of material defect detection. We provide an overview of the various types of material defects and their impact on the integrity and performance of materials. We then delve into the basic detection techniques used to identify these defects, highlighting their significance in ensuring product quality and safety. Following this, we discuss the critical role that machine learning plays in enhancing defect detection techniques, offering improvements in accuracy, efficiency, and scalability. Finally, we identify and analyze the specific challenges that current material defect detection methods face, emphasizing the need for advanced machine learning solutions to address these issues effectively.
Table 1. Common Types of Material Defects
Defect Type Description
Cracks (golewski2021validation, ) Linear separations or fractures in the material, often due to stress or impact.
Porosity (wang2022role, ) Presence of voids or pores within the material, often resulting from trapped gases during processing.
Inclusions (gong2020deep, ) Foreign materials embedded within the host material, which can compromise strength and integrity.
Corrosion (li2020point, ) Chemical or electrochemical degradation of materials, often leading to surface deterioration or structural weakness.
Delamination (segers2020backside, ) Separation of layers in laminated materials, often due to weak adhesive bonds or external stresses.
Discolouration (song2023x, ) Unintended changes in material colour, often indicative of chemical changes or degradation.
Warping (saluja2020closed, ) Deformation of the material from its intended shape, often due to uneven heat or stress distributions.
Wear (zhang2022study, ) Deterioration or degradation of materials’ physical properties over time, often as a result of repeated use, friction, or exposure to environmental factors.
2.1. Types of Material Defects
The integrity and performance of materials are pivotal in determining the reliability and longevity of engineered products and structures. A critical aspect of material science and engineering is the identification, understanding, and mitigation of material defects, which can significantly impact product quality and safety (arandiyan2021defect, ; choudhary2022recent, ). Material defects, often microscopic in nature, can escalate into macroscopic failures, leading to catastrophic outcomes in high-stakes applications such as aerospace, automotive, and civil infrastructure. The different types of material defects are listed and described in Table 1. Understanding these defects is crucial for developing strategies to detect, mitigate, or prevent them, thereby enhancing material performance and extending the lifespan of products and structures.
@@ 2.2. Defect Detection
In the manufacturing process of industrial products, various factors such as product design, processing environment, and production equipment failures often lead to the occurrence of diverse defects in the produced items, such as depressions, scratches, and voids. These defects contribute to a decline in the yield of high-quality products, escalate production costs for enterprises, and result in wastage of resources and economic losses. Importantly, defective products, once introduced into the market, pose a potential risk of causing significant safety incidents (kumbhar2023deepinspect, ; selot2023machine, ). Consequently, defect detection holds crucial value and significance in enhancing product quality and improving production efficiency, garnering joint attention from both the industrial and academic communities.
@@
Defect detection technology primarily focuses on inspecting surface defects of test samples, such as spots, pits, scratches, colour variations, and deficiencies, and obtains relevant information about the defects, including their location, type, depth, size, contours, and other characteristic features (ren2022state, ). Previous defect detection heavily relied on manual methods, exhibiting drawbacks such as high costs, low efficiency, suboptimal accuracy, and the inability to conduct large-scale inspections. However, with the maturity of information perception sensor technologies, such as visual imaging techniques, the advancement of defect detection technologies based on machine learning has been propelled (prunella2023deep, ; usamentiaga2022automated, ). This progress has significantly addressed the requirements of modern defect detection for high precision, efficiency, complexity, scalability, and real-time capabilities, thereby holding considerable research value and promising application prospects.
@@
The present study delineates the developmental trajectory of defect detection technology based on machine learning techniques over time, as illustrated in Fig. 1. Notably, the delineation is anchored around the resurgence of neural networks in 2012, demarcating the overall progression of defect detection technology into two major phases: traditional machine learning algorithms and deep learning algorithms (cao2020review, ).
Figure 1. Significant technical milestones of defect detection based on machine learning.
@@
Traditional machine learning methods include feature statistical methods, filter-based methods, and specific model methods. Defect detection methods based on feature statistics, such as Gray Level Co-occurrence Matrix (GLCM) (zhu2015yarn, ) and Local Binary Pattern (LBP) (tajeripour2007fabric, ), primarily describe defect images from a statistical perspective by analyzing the distribution of grayscale information. This approach aims to identify spatial grayscale distribution characteristics of defects for detection purposes. In addition, filter-based methods leverage signal processing principles, transforming grayscale information of defect images into the frequency domain. This allows the discovery of more easily manageable frequency domain defect features, which is particularly useful in scenarios where extracting defect features in the original grayscale domain is challenging. Specific methods include Finite Impulse Response Filter (FIR) (kumar2002defect, ), wavelet transform (li2006improving, ), Gabor filter (kumar2002defect, ), Fourier transform (chan2000fabric, ), and others. Furthermore, specific models designed for detecting surface defects in industrial products are employed, such as those based on Markov Random Field (MRF) (dogandvzic2005defect, ), Histogram of Oriented Gradients (HOG) (halfawy2014automated, ), and Deformable Parts Model (DPM) (he2019railway, ). In summary, defect detection algorithms based on traditional machine learning algorithms are well-suited for scenarios where defect features are clearly evident, defect types are limited, and there are relatively low detection requirements. These methods demonstrate a significant advantage in terms of fast detection speed.
With the resurgence of neural network technology, deep learning has experienced vibrant development and has been widely applied in defect detection, making it one of the most active research areas in the field (bhatt2021image, ). Among the various deep learning technologies relevant to defect detection, two main categories stand out. The first category involves defect detection and classification, employing techniques such as RCNN, Faster-RCNN, YOLO, SSD, etc (tulbure2022review, ; qi2020review, ). These approaches enable the detection and localization of defects in sample images, providing information on the defect’s category. The second category involves defect segmentation, utilizing algorithms like FCN, U-Net, PSPNet, DeepLab, etc (tabernik2020segmentation, ). These segmentation algorithms not only detect defects but also offer pixel-level classification and segmentation. However, compared to defect detection and classification algorithms, pixel-level segmentation algorithms are more time-consuming. It is noteworthy that defect detection solutions based on deep learning technologies can meet the majority of modern defect detection requirements. Unlike traditional machine learning algorithms that require manually designed defect features, deep learning algorithms can automatically extract defect features. Moreover, these detection algorithms demonstrate robustness in handling multiple defect types, complex backgrounds, and large-scale detection tasks. However, a drawback is that training detection models demand substantial resources, including computational power and the collection and annotation of massive training data.

### v3

@@ Abstract.
Material defects (MD) represent a primary challenge affecting product performance and giving rise to safety issues in related products. The rapid and accurate identification and localization of MD constitute crucial research endeavors in addressing contemporary challenges associated with MD. In recent years, propelled by the swift advancement of machine learning (ML) technologies, particularly exemplified by deep learning, ML has swiftly emerged as the core technology and a prominent research direction for material defect detection (MDD). Through a comprehensive review of the latest literature, we systematically survey the ML techniques applied in MDD into five categories: unsupervised learning, supervised learning, semi-supervised learning, reinforcement learning, and generative learning. We provide a detailed analysis of the main principles and techniques used, together with the advantages and potential challenges associated with these techniques. Furthermore, the survey focuses on the techniques for defect detection in composite materials, which are important types of materials enjoying increasingly wide application in various industries such as aerospace, automotive, construction, and renewable energy. Finally, the survey explores potential future directions in MDD utilizing ML technologies. This survey consolidates ML-based MDD literature and provides a foundation for future research and practice.
@@
copyright: acmcopyrightjournalyear: 2025doi: 10.1145/1122445.1122456conference: ACM Computing Surveys; ACM Computing Surveysbooktitle: ;price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Computer vision
@@ 1. Introduction
Material defect detection (MDD) stands as a critical subfield within the broader domain of defect detection, with a primary focus on identifying and assessing defects in various materials. It is integral to product quality assurance, production efficiency, and compliance with industry regulations, especially in manufacturing sectors where defects can compromise performance and safety (usamentiaga2022automated, ). MDD is a pivotal element in ensuring product reliability and safety, and it holds particular significance in the manufacturing industry. The potential impacts of defects necessitate robust detection methodologies, particularly given the increasing complexity of products and materials. Adherence to strict regulatory standards is crucial, not only to prevent catastrophic failures and reduce waste but also to maintain cost efficiency. In addition, MDD supports environmental sustainability by minimizing material waste and optimizing resource use, thus contributing to responsible production and innovation. Effective defect detection methodologies directly contribute to the longevity and performance of products, thereby driving innovation in manufacturing processes. This combination of quality assurance, safety, and sustainability underscores the critical role of MDD across diverse industrial applications.
@@
Current MDD methods must adapt to the increasing complexity of manufacturing processes and product diversity to meet the demands of modern manufacturing. The challenges in MDD arise from the need to detect intricate defects, such as cracks, pores, and inclusions, which vary in shape, size, and location. One major challenge is the identification of complex defects. Additionally, the detection of minute defects, often imperceptible to the naked eye, requires advanced high-precision methods. Furthermore, MDD often involves handling multimodal data, including images, ultrasonic waves, and X-rays. Each data type has distinct features and noise levels, posing a challenge in effectively processing and integrating these diverse sources of information (shi2020overview, ; ou2021recent, ). Timely identification and resolution of issues during production demand real-time online detection with robust performance (yang2020using, ; zeng2022small, ; zhao2023multi, ). This requirement necessitates the development of methods with fast, stable, and reliable performance. Moreover, the rarity of defective samples and the significant resources required for defect labeling create additional barriers for efficient MDD (jin2022survey, ; shahrabadi2022defect, ).
Traditional approaches to MDD primarily rely on manual inspection, which faces inherent limitations in modern industrial applications. Manual inspection depends on human expertise, making it prone to subjectivity, fatigue, and inconsistency, particularly when detecting small or subtle defects (bhatt2021image, ). The increasing complexity of manufacturing processes and products further exposes the shortcomings of manual methods, as they often fail to deliver the precision and reliability required for identifying intricate or hidden defects. Additionally, the labor-intensive nature of manual inspection limits its scalability, making it impractical for high-speed or large-scale production lines. These constraints not only hinder the efficiency of quality control processes but also increase the likelihood of undetected defects, potentially compromising product performance and safety. To address these challenges, the industrial sector has increasingly turned to automated methods, which offer enhanced accuracy, speed, and scalability. This shift has paved the way for the adoption of machine vision and other advanced technologies in MDD, laying a foundation for more reliable defect detection systems.
@@
As a response to the limitations of manual inspection, visual perception-based surface MDD has emerged as a highly effective alternative, progressively replacing traditional methods. Leveraging advancements in machine vision, these technologies enable non-contact, automated detection, providing higher accuracy and reliable performance even in complex production environments (hafiz2021attention, ; ren2022state, ). By integrating advanced image processing algorithms, visual perception-based systems are capable of automatically identifying and classifying defects, efficiently handling large data volumes, and performing high-precision inspections at industrial scale. These systems significantly enhance defect detection accuracy and consistency, minimizing reliance on human inspection and reducing the impact of human error, thereby ensuring more reliable and uniform inspection outcomes. The improved detection quality contributes directly to lowering production costs and boosting overall manufacturing efficiency (ren2022state, ; luo2020automated, ; fan2020line, ). The effectiveness of these technologies has been demonstrated across various sectors, including tile quality assurance (stojanovic2001real, ), fabric defect detection (wang2021defect, ; dong2021automatic, ), and steel plate inspection (hu2020unsupervised, ; wang2018distributed, ). Additionally, significant advancements have been achieved in PCB defect detection, as evidenced by the work of Zhang et al. (zhang2021semi, ) and Mei et al. (mei2018unsupervised, ). By seamlessly integrating these capabilities, visual perception-based surface MDD systems address the challenges of modern manufacturing, setting new standards for quality assurance and operational efficiency.
Building on the advancements of vision-based MDD, machine learning (ML), a subset of artificial intelligence (AI), has further revolutionized automated MDD. ML augments the capabilities of machine vision by automating the entire detection process, from data preprocessing and feature extraction to defect identification and classification (nturambirwe2020machine, ). Since its inception in the mid-20th century, ML has evolved from foundational concepts, such as early statistical methods, to algorithms designed to handle both structured and unstructured data, driven by advancements in computational power and the increasing availability of large datasets (janiesch2021machine, ; cioffi2020artificial, ). These advancements have enabled ML models to detect even the smallest defects with high precision, addressing challenges that are difficult for vision-based methods alone, such as subtle or hidden anomalies. ML ensures consistent performance, unaffected by human fatigue, and processes large data volumes rapidly, making it ideal for high-speed production environments. Beyond real-time applications, ML supports offline quality control by analyzing historical data to uncover defect trends and patterns, informing process improvements and reducing defect rates (sresakoolchai2022railway, ; cao2020review, ). By integrating ML with vision-based technologies, modern MDD systems achieve a new level of reliability and operational efficiency, setting benchmarks for quality assurance in manufacturing.
Despite the transformative potential of ML in MDD, several limitations hinder its broader adoption in industrial settings. One key challenge is the limited generalization ability of ML models, as detection algorithms are often tailored to specific materials or defect types, making them less effective when applied to new contexts or diverse manufacturing environments (cui2021sddnet, ). Additionally, ML-based systems require high-quality visual data for optimal performance, but industrial environments frequently present challenges such as complex backgrounds, noise, and inconsistent lighting conditions. The preparation of clean, high-resolution defect data is time-consuming and resource-intensive, posing barriers for seamless implementation. Maintenance and operational costs present further obstacles, as ML-driven systems depend on specialized hardware, advanced software, and trained personnel. The substantial initial investment, along with ongoing expenses for system upgrades, repairs, and training, can be particularly burdensome for small and medium-sized enterprises with constrained budgets (caiazzo2022towards, ). Addressing these limitations is essential to fully realize the potential of ML in MDD, enabling the development of robust, adaptable, and cost-effective solutions for diverse industrial applications.
Deep learning (DL) has emerged as a transformative ML approach in MDD, addressing limitations of traditional machine vision algorithms that often require task-specific designs. By constructing advanced network structures, DL enables the automatic extraction of high-level features from images, eliminating the need for manual feature engineering and streamlining defect detection and recognition processes (hafiz2021attention, ). This capability allows DL to effectively handle the diverse and complex requirements of contemporary MDD, including automation, high detection efficiency, accuracy, and real-time performance (bhatt2021image, ). The flexibility and adaptability of DL models make them suitable for a wide range of defect types and industrial settings, significantly advancing the field. Furthermore, the proliferation of high-performance computing resources, such as GPUs, has reduced the cost and complexity of deploying DL-based systems in MDD. These advancements have made DL not only a powerful tool for enhancing defect detection accuracy but also a practical and scalable solution for modern manufacturing challenges.
In this survey, while general ML-driven MDD techniques are extensively discussed, composite materials are highlighted separately due to their unique challenges. Unlike homogeneous materials such as metals or ceramics, composites exhibit anisotropy, heterogeneity, and complex failure mechanisms, which demand specialized defect detection methodologies (P9, ; P17, ). This distinction ensures a comprehensive exploration of both general approaches and those tailored to the specific needs of composite materials.
Table 1. Comparison of our survey with related review papers on ML for Material Defect Detection
Key Dimensions (P2, ) (P3, ) (P5, ) (P6, ) (P7, ) (P8, ) (P9, ) (P10, ) Ours
General ML Techniques for MDD
Focus on DL
Specific Material Domains
Emerging Techniques (e.g., RL, SSL)
Cross-Domain Applicability
Emphasis on Industrial Deployment
Comprehensive ML Framework for MDD
@@
The existing literature on defect detection predominantly focuses on general methodologies or specific application domains (P2, ; P6, ; P7, ; P8, ; P10, ; P11, ), with some studies delving into defect detection in particular types of materials (P3, ; P5, ; P9, ; P17, ). However, a systematic review of ML-driven MDD remains largely unexplored. As shown in Table 1, our survey addresses key gaps in the existing literature by providing a comprehensive framework and emphasizing emerging techniques. Our survey presents several distinct advantages that highlight its contributions to the field:
@@ 2. Preliminaries and Problem Identification
2.1. Acronyms and Their Definitions
Our paper introduces a substantial number of technical and academic concepts. To ensure the presentation remains clear and concise, we utilize their respective acronyms throughout the paper. For the convenience of the reader, we have included a comprehensive table of acronyms, as shown in Table 2, which provides the corresponding definitions for all acronyms used in the paper.
Table 2. Definitions of acronyms used in this study
Acronym Definition Acronym Definition
AI Artificial Intelligence APS Aluminium Profile Surface
CAE Convolutional Autoencoder CCVAE Conditional Convolutional Variational Autoencoder
CMs Composite Materials CNN Convolutional Neural Network
DBN Deep Belief Network DL Deep Learning
ECT Eddy Current Testing GAN Generative Adversarial Network
GL Genetive Learning GNNs Graph Neural Networks
GFRP Glass Fiber Reinforced Plastic HOG Histogram of Oriented Gradients
HSI Hyperspectral Imaging MD Material Defect
MDD Material Defect Detection ML Machine Learning
MPT Magnetic Particle Testing NDT Non-Destructive Testing
PCA Principal Component Analysis PT Penetrant Testing
RL Reinforcement Learning RNNs Recurrent Neural Networks
SVM Support Vector Machine SL Supervised Learning
SSL Semi-Supervised Learning UL Unsupervised Learning
@@ 2.2. Defect Detection
Defect detection technology primarily focuses on inspecting surface defects of test samples, such as spots, scratches, colour variations, and deficiencies, and obtains relevant information about the defects, including their location, type, size, contours, and other characteristic features (ren2022state, ). Previous defect detection heavily relied on manual methods, exhibiting drawbacks such as high costs, low efficiency, suboptimal accuracy, and the inability to conduct large-scale inspections. However, with the maturity of information perception sensor technologies, such as visual imaging techniques, the advancement of defect detection technologies based on ML has been propelled (prunella2023deep, ; usamentiaga2022automated, ).
@@
The present study delineates the developmental trajectory of defect detection technology based on ML techniques over time, as illustrated in Fig. 1. Notably, the delineation is anchored around the resurgence of neural networks in 2012, demarcating the overall progression of defect detection technology into two major phases: traditional ML algorithms and DL algorithms (cao2020review, ).
@@
Traditional ML methods for defect detection include feature statistical methods, filter-based methods, and specific model methods. Statistical methods like Gray Level Co-occurrence Matrix (GLCM) (zhu2015yarn, ) and Local Binary Pattern (LBP) (tajeripour2007fabric, ) analyze spatial grayscale distributions to identify defects. Filter-based methods leverage signal processing to extract frequency domain features, utilizing techniques such as wavelet transform (li2006improving, ), Gabor filter (kumar2002defect, ), and Fourier transform (chan2000fabric, ). Specific models, such as Markov Random Field (MRF) (dogandvzic2005defect, ) and Histogram of Oriented Gradients (HOG) (halfawy2014automated, ), are also employed for surface defect detection. These traditional methods excel in scenarios with limited defect types and low detection complexity, offering fast detection speeds.
With the resurgence of neural networks, DL has become a prominent area in defect detection (bhatt2021image, ). Two main DL categories stand out: defect detection/classification and defect segmentation. Detection/classification techniques, such as RCNN, Faster-RCNN, YOLO, and SSD, identify and localize defects while categorizing them (tulbure2022review, ; qi2020review, ). Segmentation algorithms, like U-Net and DeepLab, enable pixel-level classification but are more time-intensive (tabernik2020segmentation, ). Unlike traditional ML methods, DL algorithms automatically extract features and handle complex backgrounds and multi-defect scenarios effectively. However, they require significant computational resources and extensive annotated data.
Figure 1. Significant technical milestones of defect detection based on ML.